OpenStudy (anonymous):

solve using substitution 3x + 2y = -4 x + 2y = -8

OpenStudy (anonymous):

just substitue in 0 to wherever u see y for the x intercept

OpenStudy (anonymous):

why 0?

OpenStudy (anonymous):

because when y is 0, you will find your x intercept

OpenStudy (ipwnbunnies):

Err, Lucy that is incorrect. We're trying to solve 2 linear equations, not one.

OpenStudy (anonymous):

okay, how do i solve it @iPwnBunnies

OpenStudy (anonymous):

3x+2y=-4 x+2y=-8 x=-2y-8 3(x)+2y=-4 ----->3(-2y-8)+2y=-4 Solve for y then plug in y in the x= equation

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

so it would be (2,-5)?

OpenStudy (anonymous):

First, solve for one of the variables in one of the equations. I'll do x in the second equation: \[x + 2y = -8 \] \[x = -8 - 2y\] Great! Now plug this x into the other equation: \[3 (-8 - 2y) + 2y = -4 \] \[-24 - 6y + 2y = -4\] \[-4y = 20\] \[y = - 5\] We can plug this result into either equation to get y: \[x + 2(-5) = -8\] \[x - 10 = -8\] \[x = 2\] We can check by plugging x and y into both equations. Hope this helps!

OpenStudy (anonymous):

I'm great at late responses, sorry about that! :P